OpenStudy (anonymous):

Can someone help me please? :) An expression is shown below: 2x^3y + 8xy - 4x^2y - 16y Part A ~ Rewrite the expression so that the GCF is factored completely. Part B ~ Rewrite the expression completely factored. Show the steps of your work. Part C ~ If the two middle terms were switched so that the expression became 2x^3y - 4x^2y + 8xy - 16y, would the factored expression no longer be equivalent to your answer on Part B? Explain your reasoning.

OpenStudy (texaschic101):

2y(x^3 + 4x - 2x^2 - 8) ==== A 2y(x - 2)(x^2 + 4) ===== B if 2 middle terms were switched.. 2x^3y - 4x^2y + 8xy - 16y 2y(x^3 - 2x^2 + 4x - 8) 2y(x^2 + 4)(x - 2) ==== It is equivalent to answer on part B

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

Do you think you can show how you got step B please?

OpenStudy (texaschic101):

I don't really know how to explain it....it is kinda like the reverse of the FOIL method.

OpenStudy (texaschic101):

@eliassaab ..can you help with this. I do not really know how to explain to HMT how I got step B

OpenStudy (anonymous):

It is the commutative property of addition - 4x^2y + 8xy = + 8xy - 4x^2y So nothing change. The two quantities are the same.

OpenStudy (texaschic101):

I think he is asking how I completely factored..not just the GCF, but the reverse FOIL stuff

OpenStudy (anonymous):

Ok, first I'm a she. And second, I know that changing the order doesn't make a difference in the answer. I was wondering if you guys could show me the steps of how you the answer in part B.

OpenStudy (anonymous):

* got the answer

OpenStudy (anonymous):

Two equal expression will have the same factorization. It is a simple as that.

OpenStudy (texaschic101):

sorry bout that....she

OpenStudy (anonymous):

It's all good.

OpenStudy (texaschic101):

Do you know the FOIL method ?

OpenStudy (anonymous):

F ~ First term * first term I ~ Inside term * inside term O ~ Outside term * outside term L ~ Last term * last term

OpenStudy (texaschic101):

ok...good..now look at this..2y(x^3 + 4x - 2x^2 - 8) with FOIL ... 2y (x - 2)(x^2 + 4) first ... x * x^2 = 3x^2 outside .. x * 4 = 4x inside .. -2 * x^2 = -2x^2 last .. -2 * 4 = 8 when they are put all together, they equal 3x^2 + 4x - 2x^2 + 8....it equals. I am not sure how to explain how to do this....I do it in my head.

OpenStudy (texaschic101):

I am sorry that I can't help more. You might want to post another question relating to how to do the reverse FOIL and somebody else can probably explain it better to you

OpenStudy (anonymous):

Ohh, ok. I got it! No, that's ok. I just needed basic instructions, the teachers teach it all crazy and it gets confusing. Yours was much easier to understand. Thank you soo much!

OpenStudy (texaschic101):

some teachers are like that...lol.

OpenStudy (anonymous):

Haha yea.